OpenStudy (anonymous):

Ms. Rao buys a computer, a printer, and a scanner for $2,543. The computer cost $1,502 more then the printer. The printer cost $123 more than the scanner. How much does the computer cost? can someone please help me?

OpenStudy (anonymous):

let the price of the computer be 'c' the price of printer be 'p' and the price of the scanner be 's' then p + c + s = 2543 and c = p + 1503 p = 123 + s three equations, three variable should be simple enough from here.

OpenStudy (anonymous):

thanks but am i figuring out the cost of the scanner or printer first?

OpenStudy (anonymous):

yeah. you have to.

OpenStudy (anonymous):

is my answer 1808?

OpenStudy (anonymous):

I guess its 1890.66

OpenStudy (anonymous):

ok i see i typed wrong. It should be 1502. not 3. my mistake.